# SpargeAttn: Accurate Sparse Attention Accelerating Any Model Inference

### v2

@@ Abstract
An efficient attention implementation is essential for large models due to its quadratic time complexity. Fortunately, attention commonly exhibits sparsity, i.e., many values in the attention map are near zero, allowing for the omission of corresponding computations. Many studies have utilized the sparse pattern to accelerate attention. However, most existing works focus on optimizing attention within specific models by exploiting certain sparse patterns of the attention map. A universal sparse attention that guarantees both the speedup and end-to-end performance of diverse models remains elusive. In this paper, we propose SpargeAttn, a universal sparse and quantized attention for any model. Our method uses a two-stage online filter: in the first stage, we rapidly and accurately predict the attention map, enabling the skip of some matrix multiplications in attention. In the second stage, we design an online softmax-aware filter that incurs no extra overhead and further skips some matrix multiplications. Experiments show that our method significantly accelerates diverse models, including language, image, and video generation, without sacrificing end-to-end metrics. The codes are available at https://github.com/thu-ml/SpargeAttn.  [^1] All experiments using SpargeAttn is based on SageAttention. An updated implementation based on SageAttention2, is available at https://github.com/thu-ml/SpargeAttn. It further offers a 30% speedup over the Attention in this paper.
@@ 1 Introduction
As sequence lengths in large models become longer, such as 45K-128K in video generation and language models (Yang et al., 2025; Bao et al., 2024; Dubey et al., 2024), the time consuming of attention occupies a significant portion of inference latency in large models (Zhang et al., 2025). Fortunately, the attention map $P=\mathrm{Softmax}(QK^{\top}/\sqrt{d})$ exhibits inherent sparsity, as the softmax operation often creates many values approaching zero (Deng et al., 2024). Sparse attention methods exploit such sparsity to accelerate attention by (1) constructing a “sparse mask”, which indicates the important non-zero entries of the attention map $P$ that should be computed, and (2) computing attention only for the parts corresponding to the sparse mask. There are three distinct categories of sparse attention methods based on how the sparse mask is generated. pattern-based method (Zhang et al., 2023; Xiao et al., 2024a; Fu et al., 2024; Zhu et al., 2024; Xiao et al., 2025, 2024b) relies on specific sparsity patterns based on empirical observations, dynamic sparse attention (Ribar et al., 2024; Singhania et al., 2024; Jiang et al., 2024; FlexPrefill, 2025; Gao et al., 2024) computes the mask on-the-fly based on the inputs, and training-based method (Kitaev et al., 2020; Pagliardini et al., 2023) directly train models with native sparse attention.
@@
Our approach. In this work, we develop SpargeAttn, a training-free sparse attention that can be adopted universally on various tasks, including language modeling and text-to-image/video, and various sequence lengths. We propose three main techniques to improve the universality, accuracy, and efficiency. First, we propose a universal sparse mask prediction algorithm, which constructs the sparse mask by compressing each block of $Q$, $K$ to a single token. Importantly, we compress selectively based on the similarity of tokens within the block, so the algorithm can accurately predict sparse masks universally across tasks. Second, we propose a sparse online softmax algorithm at the GPU warp level, which further omits some $PV$ products by leveraging the difference between global maximum values and local maximum values in online softmax. Third, we integrate this sparse approach into the 8-bit quantized SageAttention framework for further acceleration.
@@ 2 Related Work
Depending on how the sparsity mask is constructed, sparse attention methods can be divided into three types: (1) Pattern required methods rely on some fixed patterns of the attention map, such as sliding windows or attention sinks (Xiao et al., 2024b). H2O (Zhang et al., 2023), InfLLM (Xiao et al., 2024a), and DUOAttention (Xiao et al., 2025) rely on sliding window pattern. SampleAttention (Zhu et al., 2024), MOA (Fu et al., 2024), and StreamingLLM (Xiao et al., 2024b) rely on sliding window and attention sink pattern. DitFastAttn (Yuan et al., 2024) relies on sliding window patterns and similarities between different attention maps. Moreover, DitFastAttn is restricted to simple diffusion transformers, showing incompatibility with language models and MMDiT models like Flux (Black Forest Labs, 2023), Stable Diffusion3 and 3.5 (Stability AI, 2023), and CogVideoX (Yang et al., 2025). As the pattern varies across models, these methods may not universally work for different models. (2) Dynamic sparse methods dynamically construct the sparse mask based on the input without the need of preset patterns, and are thus potentially more universal. Existing works can be further categorized into channel compression and token compression. Channel compression methods include SparQAttn (Ribar et al., 2024) and LokiAttn (Singhania et al., 2024). They construct the mask by carrying full attention with reduced dimensionality. However, as the dimension is already small, e.g., 64, 128, in commonly used attention, the speedup potential might be limited. Token compression methods include MInference (Jiang et al., 2024) and FlexPrefill (FlexPrefill, 2025). They construct the mask by compressing each block of tokens to a single token and compute attention on this shorter sequence. However, this approximation is too aggressive: missing important blocks of $P$ is possible if they do not have a large attention score on the compressed sequence. SeerAttention (Gao et al., 2024) requires training of additional parameters for attention, which is expensive to use. Moreover, they are all designed for language models, and their applicability to other model types, such as diffusion models, remains uncertain. (3) Training-based methods modify the attention computation logic, requiring retraining the entire model, such as Reformer (Kitaev et al., 2020) and FastAttention (Pagliardini et al., 2023). These methods are much more expensive to use than training-free methods.
There are other ways to accelerate attention, such as optimizing the kernel implementation (Dao et al., 2022; Dao, 2024; Shah et al., 2024), quantization (Zhang et al., 2025), distributing the workload (Liu et al., 2024a), and designing linear time attention (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020). They are orthogonal to our approach.
@@ 3.1 Sparse FlashAttention
where $m_{ij}$ and $l_{ij}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively. The $\tilde{\sigma}()$ is an operator similar to softmax.: $m_{ij}=\max\{m_{i,j-1},\mathrm{rowmax}(S_{ij})\},~{}\widetilde{P}_{i,j}=\exp(S_{ij}-m_{ij})$. Finally, the output $O_{i}$ can be computed by $O_{i}=\mathrm{diag}(l_{ij})^{-1}O_{ij}$.
@@
20:    $O_{i}=\mathrm{diag}(l_{i,T_{n}})^{-1}O_{i,T_{n}}$ ;    Write $O_{i}$ ;
21:  end for
22:  return $O=\{O_{i}\}$ ;
@@ 3.2 Selective Token Compression for Sparse Prediction
Key idea. Although attention maps vary across models, we observe that various models exhibit a common trait: Most closer tokens in the query and key matrices of the attention show high similarity (See Fig. 4). Consequently, for blocks composed of highly similar tokens, we can consolidate these tokens into a single representative token for the block. Based on this observation, we propose a pattern-free online prediction method for identifying sparse blocks in $P$ to skip some computation of $Q_{i}K_{j}^{\top}$ and $\widetilde{P}_{ij}V_{j}$ during the FlashAttention process. Specifically, we first compress blocks exhibiting high self-similarity within $Q$ and $K$ into tokens. Then, we swiftly compute a compressed attention map $\hat{P}$ using the compressed $Q$ and $K$. Finally, we selectively compute $\{Q_{i}K_{j}^{\top},\widetilde{P}_{ij}V_{j}\}$ for those pairs $(i,j)$ where $\{\hat{P}[i,j]\}$ accumulates a high score in the compressed attention map. Importantly, compressing only the token blocks with high self-similarity is crucial, as omitting computations for non-self-similar blocks can result in the loss of critical information. This will be confirmed in Sec. 4 and A.2.
@@
where $Q_{i}\in\mathbb{R}^{b_{q}\times d},q_{i}\in\mathbb{R}^{1\times d},K_{j}\in\mathbb{R}^{b_{k}\times d},k_{j}\in\mathbb{R}^{1\times d}$ and $\mathrm{CosSim}(X)=\frac{XX^{\top}}{|\max(XX^{\top})|}$ measures the cosine-similarity within a block.
@@
Finally, we need to ensure that calculations involving non-self-similar blocks of $Q$ or $K$ are not omitted. Therefore, we set all values in the rows of $M_{g}$ corresponding to not self-similar blocks of $Q$ to 1, and all values in the columns of $M_{g}$ corresponding to non-self-similar blocks of $K$ to 1.
@@ 3.3 Masking of the First Stage
Masking. The $M_{g}$ can be applied in FlashAttention directly to saving some computation. In the inner loop of FlashAttention, i.e., during computing attention between a $Q_{i}$ and $\{K_{j}\},\{V_{j}\}$, we can skip {$Q_{i}K_{j}^{\top}$, $\widetilde{P}_{ij}V_{j}$} when $M_{g}[i,j]=0$.
@@ 3.4 Sparse Warp Online Softmax
If $\mathrm{rowmax}(S_{ij})<m_{ij}$, then $m_{ij}=m_{i,j-1}$. Consequently, $O_{ij}=O_{i,j-1}+\widetilde{P}_{ij}V_{j}$. Furthermore, if $\mathrm{rowmax}(S_{ij})\ll m_{ij}$ holds ture, then all values in $\widetilde{P}_{ij}=\exp(S_{ij}-m_{ij})$ are close to 0. This results in all values in $\widetilde{P}_{ij}V_{j}$ being close to 0. This condition implies that $\widetilde{P}_{ij}V_{j}$ is negligible when $\mathrm{rowmax}(S_{ij})$ is significantly smaller than $m_{ij}$:
@@ 3.5 Combined with SageAttention
To further accelerate our implementation of sparse attention, we integrate our method into SageAttention (Zhang et al., 2025), which proposes a quantized method for accelerating attention. Since quantization operations and sparse operations are orthogonal, sparse computation can be directly applied to SageAttention. The complete algorithm is shown in Algorithm 1. Specifically, first, we need to add one judgment at the beginning of the inner loop of SageAttention (Line 10 in Algorithm 1) to decide whether to skip the whole inner loop once. Second, we add another judgment before the updating of $O_{ij}$ in the inner loop of SageAttention (Line 15 in Algorithm 1) to decide whether to skip the computation of $\widetilde{P}_{ij}V_{j}$. Moreover, to minimize the attention map prediction overhead, we implement the prediction using CUDA and adopt some kernel fusion techniques.
@@ 3.7 HilbertCurve Permutation
Key idea. Improving sparsity while maintaining accuracy is a key challenge in enhancing the performance of sparse attention. In our algorithm, increasing the self-similarity of key and query blocks can reduce the number of non-self-similar blocks. This allows more blocks to participate in $\mathrm{TopCdf}$ selection, thereby improving sparsity. Since attention is computationally invariant to token permutations, the problem reduces to finding a permutation that enhances the similarity of adjacent tokens.
Image and video models benefit from strong priors: adjacent pixels are likely to be similar. To better leverage this prior, we propose the HilbertCurve permutation, given 3D visual tokens $Q,K,V\in\mathbb{R}^{T\times H\times W\times d}$, We use the Hilbert Curve to fill the 3D space and then flatten tokens along the curve into shape $\mathbb{R}^{L\times d},L=T\times H\times W$. Fig. 5 illustrates an example of $1\times 6\times 6$ visual tokens flatten by row-major order and HilbertCurve. The Hilbert Curve preserves locality effectively, traversing the entire 3D space without crossing rows or columns, thereby increasing the similarity of adjacent tokens and the sparsity of attention.
Table 1: End-to-end metrics across text, image, and video generation models. ✗ indicates an inability to generate results for evaluation. The speed and sparsity are the average for each layer in the model in real generation tasks described in Sec. 4.1. The speed and sparsity of Llama3.1 are measured in the NeedleInAHaystack task with a 128K sequence length.
@@ 4.1 Setup
Models. We validate the effectiveness of SpargeAttn across diverse representative models from language, image, and video generation. Specifically, we conduct experiments on Llama3.1 (8B) (Dubey et al., 2024) for text-to-text, CogvideoX (2B) and Mochi (Team, 2024) for text-to-video, Flux (Black Forest Labs, 2023)(.1-dev) and Stable-Diffusion3.5 (large) (Stability AI, 2023) for text-to-image.
@@
Datasets. The Text-to-text model is evaluated on four zero-shot tasks: WikiText (Merity et al., 2017) to assess the model’s prediction confidence, Longbench (Bai et al., 2024) and En.MC of InfiniteBench (Zhang et al., 2024b) for a comprehensive assessment of long context understanding capabilities, and the Needle-in-A-Haystack task (Kamradt, 2023) to assess the model’s retrieval ability. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets. Text-to-image models are assessed on COCO annotations (Lin et al., 2014).
@@
End-to-end metrics. For Llama3.1, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Longbench score (Bai et al., 2024), and retrival accuracy for the Needle-in-A-Haystack task (Kamradt, 2023). For text-to-video models, following Zhao et al. (2025), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024b) to measure the text-video alignment; VQA-a and VQA-t to assess the video aesthetic and technical quality, and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in the COCO dataset in three aspects: FID (Heusel et al., 2017) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference.
@@
Speed and sparsity metric. We use TOPS (tera operations per second) to evaluate the speed of sparse attention methods. Specifically, TOPS = $O(attn)/t$, where $O(attn)$ represents the total number of operations in a standard attention computation, and $t$ is the latency from a given $(Q,K,V)$ to the output of attention. Note that this speed metric is completely fair. This is because the $O(attn)$ is fixed for a set of inputs, and then the speed is determined by $t$, which includes the time spent predicting the sparse region of the attention map. We define Sparsity as the proportion of the Matmul of $Q_{i}K_{j}$ plus $P_{i}^{j}V_{j}$ that are skipped relative to the total number of $Q_{i}K_{j}$ plus $P_{i}^{j}V_{j}$ in a full attention required.
@@
Implementation and Hyper-parameters. We implement our method using CUDA. As discussed in Sec. 3.6, we need to determine $l1,l2$ for models. We use ($l_{1}=0.08,l_{2}=0.09$) for Llama3.1, ($l_{1}=0.05,l_{2}=0.06$) for CogvideoX and Mochi, and ($l_{1}=0.07,l_{2}=0.08$) for Stable-Diffusion3.5 and Flux.
@@
Baselines. Currently, sparse attention methods applicable across different model types are limited. We choose block-sparse MInference (Jiang et al., 2024) and FlexPrefill (FlexPrefill, 2025) as our baselines. To vary the sparsity of these baselines, we use 30% and 70% for MInference, and use $\gamma=0.95$ and $0.99$ for FlexPrefill according to their paper.
@@
Figure 8: A NeedleInAHaystack comparison example on Llama3.1. The sparsity of SpargeAttn, MInference, and FlexPrefill is 0.5, 0.5, and 0.54.

### v4

# SpargeAttention: Accurate and Training-free Sparse Attention Accelerating Any Model Inference

@@ Abstract
An efficient attention implementation is essential for large models due to its quadratic time complexity. Fortunately, attention commonly exhibits sparsity, i.e., many values in the attention map are near zero, allowing for the omission of corresponding computations. Many studies have utilized the sparse pattern to accelerate attention. However, most existing works focus on optimizing attention within specific models by exploiting certain sparse patterns of the attention map. A universal sparse attention that guarantees both the speedup and end-to-end performance of diverse models remains elusive. In this paper, we propose SpargeAttn, a universal sparse and quantized attention for any model. Our method uses a two-stage online filter: in the first stage, we rapidly and accurately predict the attention map, enabling the skip of some matrix multiplications in attention. In the second stage, we design an online softmax-aware filter that incurs no extra overhead and further skips some matrix multiplications. Experiments show that our method significantly accelerates diverse models, including language, image, and video generation, without sacrificing end-to-end metrics. The code is available at https://github.com/thu-ml/SpargeAttn.  [^1] All experiments using SpargeAttn is based on SageAttention. An updated implementation based on SageAttention2, is available at https://github.com/thu-ml/SpargeAttn. It further offers a 30% speedup over the attention in this paper.
@@ 1 Introduction
As sequence lengths in large models become longer, such as 45K-128K in video generation and language models (Yang et al., 2025b; Bao et al., 2024; Dubey et al., 2024; Zhang et al., 2025c), the time consuming of attention occupies a significant portion of inference latency in large models (Zhang et al., 2025d). Fortunately, the attention map $P=\mathrm{Softmax}(QK^{\top}/\sqrt{d})$ exhibits inherent sparsity, as the softmax operation often creates many values approaching zero (Deng et al., 2024). Sparse attention exploit such sparsity to accelerate attention by (1) constructing a “sparse mask”, which indicates the important non-zero entries of the attention map $P$ that should be computed, and (2) computing attention only for the parts corresponding to the sparse mask. There are three distinct categories of sparse attention methods based on how the sparse mask is generated. pattern-based method (Zhang et al., 2023; Xiao et al., 2024a; Fu et al., 2024; Zhu et al., 2024; Xiao et al., 2025, 2024b; Zhang et al., 2025h) relies on specific sparsity patterns based on empirical observations, dynamic sparse attention (Ribar et al., 2024; Singhania et al., 2024; Jiang et al., 2024; Lai et al., 2025; Gao et al., 2024; Xi et al., 2025; Yang et al., 2025a; Zhang et al., 2025f) computes the mask on-the-fly based on the inputs, and training-based method (Kitaev et al., 2020; Pagliardini et al., 2023; Zhang et al., 2025i) directly train models with native sparse attention.
@@
Our approach. In this work, we develop SpargeAttn, a training-free sparse attention that can be adopted universally on various tasks, including language modeling and text-to-image/video, and various sequence lengths. We propose three main techniques to improve the universality, accuracy, and efficiency. First, we propose a universal sparse mask prediction algorithm, which constructs the sparse mask by compressing each block of $Q$, $K$ to a single token. Importantly, we compress selectively based on the similarity of tokens within the block, so the algorithm can accurately predict sparse masks universally across tasks. Second, we propose a sparse online softmax algorithm at the GPU warp level, which further omits some $PV$ products by leveraging the difference between global maximum values and local maximum values in online softmax. Third, we integrate this sparse approach into the 8-bit quantized SageAttention framework for further acceleration. To the best of our knowledge, SpargeAttn is the first sparse attention method that can actually accelerate across language, image, and video models without compromising accuracy.
@@ 2 Related Work
Depending on how the sparsity mask is constructed, sparse attention methods can be divided into three types: (1) Pattern required methods rely on some fixed patterns of the attention map, such as sliding windows or attention sinks (Xiao et al., 2024b). H2O (Zhang et al., 2023), InfLLM (Xiao et al., 2024a), and DUOAttention (Xiao et al., 2025) rely on sliding window pattern. SampleAttention (Zhu et al., 2024), MOA (Fu et al., 2024), and StreamingLLM (Xiao et al., 2024b) rely on sliding window and attention sink pattern. DitFastAttn (Yuan et al., 2024) relies on sliding window patterns and similarities between different attention maps. Moreover, DitFastAttn is restricted to simple diffusion transformers, showing incompatibility with language models and MMDiT models like Flux (Black Forest Labs, 2023), Stable Diffusion3 and 3.5 (Stability AI, 2023), and CogVideoX (Yang et al., 2025b). As the pattern varies across models, these methods may not universally work for different models. (2) Dynamic sparse methods dynamically construct the sparse mask based on the input without the need of preset patterns, and are thus potentially more universal. Existing works can be further categorized into channel compression and token compression. Channel compression methods include SparQAttn (Ribar et al., 2024) and LokiAttn (Singhania et al., 2024). They construct the mask by carrying full attention with reduced dimensionality. However, as the dimension is already small, e.g., 64, 128, in commonly used attention, the speedup potential might be limited. Token compression methods include MInference (Jiang et al., 2024) and FlexPrefill (Lai et al., 2025). They construct the mask by compressing each block of tokens to a single token and compute attention on this shorter sequence. However, this approximation is too aggressive: missing important blocks of $P$ is possible if they do not have a large attention score on the compressed sequence. SeerAttention (Gao et al., 2024) requires training of additional parameters for attention, which is expensive to use. Moreover, they are all designed for language models, and their applicability to other model types, such as diffusion models, remains uncertain. (3) Training-based methods modify the attention computation logic, requiring retraining the entire model, such as Reformer (Kitaev et al., 2020) and FastAttention (Pagliardini et al., 2023). These methods are much more expensive to use than training-free methods.
There are other ways to accelerate attention, such as optimizing the kernel implementation (Dao et al., 2022; Dao, 2024; Shah et al., 2024), quantization (Zhang et al., 2025d), distributing the workload (Liu et al., 2024a), and designing linear time attention (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020). They are orthogonal to our approach.
@@ 3.1 Sparse FlashAttention
where $m_{ij}$ and $l_{ij}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively. The $\tilde{\sigma}()$ is an operator similar to softmax.: $m_{ij}=\max\{m_{i,j-1},\mathrm{rowmax}(S_{ij})\},~{}\widetilde{P}_{ij}=\exp(S_{ij}-m_{ij})$. Finally, the output $O_{i}$ can be computed by $O_{i}=\mathrm{diag}(l_{ij})^{-1}O_{ij}$.
@@
20:    $O_{i}=\mathrm{diag}(l_{i,T_{n}})^{-1}O_{i,T_{n}}$ ;
21:    Write $O_{i}$ ;
22:  end for
23:  return $O=\{O_{i}\}$ ;
@@ 3.2 Selective Token Compression for Sparse Prediction
Key idea. Although attention maps vary across models, we observe that various models exhibit a common trait: Most neighboring tokens in the query and key matrices of the attention show high similarity (See Fig. 4). Consequently, for blocks composed of highly similar tokens, we can consolidate these tokens into a single representative token for the block. Based on this observation, we propose a pattern-free online prediction method for identifying sparse blocks in $P$ to skip some computation of $Q_{i}K_{j}^{\top}$ and $\widetilde{P}_{ij}V_{j}$ during the FlashAttention process. Specifically, we first compress blocks exhibiting high self-similarity within $Q$ and $K$ into tokens. Then, we swiftly compute a compressed attention map $\hat{P}$ using the compressed $Q$ and $K$. Finally, we selectively compute $\{Q_{i}K_{j}^{\top},\widetilde{P}_{ij}V_{j}\}$ for those pairs $(i,j)$ where $\{\hat{P}[i,j]\}$ accumulates a high score in the compressed attention map. Notably, block selection was only performed in the high self-similarity blocks, which we also refer to as ”selective blocks.” For those non-self-similar blocks, as a good presentation token for the whole block is hard to find, we choose to always compute the non-self-similar block in the attention operation, which we also refer to as ”fix blocks.” Importantly, compressing only the token blocks with high self-similarity is crucial, as omitting computations for fix blocks can result in the loss of critical information. This will be confirmed in Sec. 4 and A.2.
@@
where $Q_{i}\in\mathbb{R}^{b_{q}\times d},q_{i}\in\mathbb{R}^{1\times d},K_{j}\in\mathbb{R}^{b_{k}\times d},k_{j}\in\mathbb{R}^{1\times d}$ and $\mathrm{CosSim}(X)=mean(\frac{XX^{\top}}{|\max(XX^{\top})|})$ measures the cosine-similarity within a block.
@@
Finally, we need to ensure that calculations involving non-self-similar blocks(fix block) of $Q$ or $K$ are not omitted. Therefore, we set all values in the rows of $M_{g}$ corresponding to not self-similar blocks of $Q$ to 1, and all values in the columns of $M_{g}$ corresponding to non-self-similar blocks of $K$ to 1.
@@ 3.3 Masking of the First Stage
Masking. The $M_{g}$ can be applied in FlashAttention directly to save some computation. In the inner loop of FlashAttention, i.e., during computing attention between a $Q_{i}$ and $\{K_{j}\},\{V_{j}\}$, we can skip {$Q_{i}K_{j}^{\top}$, $\widetilde{P}_{ij}V_{j}$} when $M_{g}[i,j]=0$.
@@ 3.4 Sparse Warp Online Softmax
If $\mathrm{rowmax}(S_{ij})<m_{ij}$, then $m_{ij}=m_{i,j-1}$. Consequently, $O_{ij}=O_{i,j-1}+\widetilde{P}_{ij}V_{j}$. Furthermore, if $\mathrm{rowmax}(S_{ij})\ll m_{ij}$ holds true, then all values in $\widetilde{P}_{ij}=\exp(S_{ij}-m_{ij})$ are close to 0. This results in all values in $\widetilde{P}_{ij}V_{j}$ being close to 0. This condition implies that $\widetilde{P}_{ij}V_{j}$ is negligible when $\mathrm{rowmax}(S_{ij})$ is significantly smaller than $m_{ij}$:
@@ 3.5 Combined with SageAttention
To further accelerate our implementation of sparse attention, we integrate our method into SageAttention (Zhang et al., 2025a, d, g, b, e), which proposes a quantized method for accelerating attention. Since quantization (Hu et al., 2025; Zhang et al., 2025j) operations and sparse operations are orthogonal, sparse computation can be directly applied to SageAttention. The complete algorithm is shown in Algorithm 1. Specifically, first, we need to add one judgment at the beginning of the inner loop of SageAttention (Line 10, Algorithm 1) to decide whether to skip the whole inner loop once. Second, we add another judgment before the updating of $O_{ij}$ in the inner loop of SageAttention (Line, in Algorithm 1) to decide whether to skip the computation of $\widetilde{P}_{ij}V_{j}$. Moreover, to minimize the attention map prediction overhead, we implement the prediction using CUDA and adopt some kernel fusion techniques.
@@ 3.7 HilbertCurve Permutation
Key idea. Improving sparsity while maintaining accuracy is a key challenge in enhancing the performance of sparse attention. In our algorithm, increasing the self-similarity of key and query blocks can reduce the number of fix blocks. This allows more selective blocks to participate in $\mathrm{TopCdf}$ selection, thereby improving sparsity. Since attention is computationally invariant to token permutations, the problem reduces to finding a permutation that enhances the similarity of adjacent tokens.
Image and video models benefit from strong priors: adjacent pixels are likely to be similar. To better leverage this prior, we propose the HilbertCurve permutation, given 3D visual tokens $Q,K,V\in\mathbb{R}^{T\times H\times W\times d}$, We use the Hilbert Curve to fill the 3D space and then flatten tokens along the curve into shape $\mathbb{R}^{L\times d},L=T\times H\times W$. Fig. 5 illustrates an example of $1\times 6\times 6$ visual tokens flattened by row-major order and HilbertCurve. The Hilbert Curve preserves locality effectively, traversing the entire 3D space without crossing rows or columns, thereby increasing the similarity of adjacent tokens and the sparsity of attention.
Table 1: End-to-end metrics across text, image, and video generation models. ✗ indicates an inability to generate results for evaluation. The speed and sparsity are the average for each layer in the model in real generation tasks described in Sec. 4.1. The speed and sparsity of Llama3.1 are measured in the Needle-in-a-Haystack task with a 128K sequence length.
@@
Model (seq_len) Attention (Sparsity) CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$ Latency $\downarrow$
Open-Sora-Plan (38K) Full-Attention 0.1650 0.9994 81.40 80.60 0.847 629s
SpargeAttn (0.34) 0.1686 0.9985 77.59 76.91 0.839 393s
@@ 4.1 Setup
Models. We validate the effectiveness of SpargeAttn across diverse representative models from language, image, and video generation. Specifically, we conduct experiments on Llama3.1 (8B) (Dubey et al., 2024) for text-to-text, CogvideoX (2B), Mochi (Team, 2024), and Open-Sora-Plan (Lin et al., 2024) for text-to-video, Flux (.1-dev) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (large) (Stability AI, 2023) for text-to-image.
@@
Datasets. The Text-to-text model is evaluated on four zero-shot tasks: WikiText (Merity et al., 2017) to assess the model’s prediction confidence, Longbench (Bai et al., 2024) and En.MC of InfiniteBench (Zhang et al., 2024) for a comprehensive assessment of long context understanding capabilities, and the Needle-in-a-Haystack task (Kamradt, 2023) to assess the model’s retrieval ability. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024c) prompt sets. Text-to-image models are assessed on COCO annotations (Lin et al., 2014).
@@
End-to-end metrics. For Llama3.1, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Longbench score (Bai et al., 2024), and retrival accuracy for the Needle-in-a-Haystack task (Kamradt, 2023). For text-to-video models, following Zhao et al. (2025b), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024b) to measure the text-video alignment; VQA-a and VQA-t to assess the video aesthetic and technical quality, and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in the COCO dataset in three aspects: FID (Heusel et al., 2017) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference.
@@
Speed and sparsity metric. We use TOPS (tera operations per second) to evaluate the speed of sparse attention methods. Specifically, TOPS = $O(attn)/t$, where $O(attn)$ represents the total number of operations in a standard attention computation, and $t$ is the latency from a given $(Q,K,V)$ to the output of attention. Note that this speed metric is completely fair. This is because the $O(attn)$ is fixed for a set of inputs, and then the speed is determined by $t$, which includes the time spent predicting the sparse region of the attention map. We define Sparsity as the proportion of the Matmul of $Q_{i}K_{j}^{\top}$ plus $\widetilde{P}_{ij}V_{j}$ that are skipped relative to the total number of $Q_{i}K_{j}^{\top}$ plus $\widetilde{P}_{ij}V_{j}$ in a full attention required.
@@
Implementation and Hyper-parameters. We implement our method using CUDA. As discussed in Sec. 3.6, we need to determine $l_{1},l_{2}$ for models. We use ($l_{1}=0.08,l_{2}=0.09$) for Llama3.1, ($l_{1}=0.05,l_{2}=0.06$) for CogvideoX and Mochi, and ($l_{1}=0.07,l_{2}=0.08$) for Stable-Diffusion3.5 and Flux, ($l_{1}=0.03,l_{2}=0.035$) for Open-Sora-Plan.
@@
Baselines. Currently, sparse attention methods applicable across different model types are limited. We choose block-sparse MInference (Jiang et al., 2024) and FlexPrefill (Lai et al., 2025) as our baselines. To vary the sparsity of these baselines, we use 30% and 70% for MInference, and use $\gamma=0.95$ and $0.99$ for FlexPrefill according to their paper.
Figure 6: Visible examples on CogvideoX using SpargeAttention.
@@
Figure 9: A Needle-in-a-Haystack comparison example on Llama3.1. The sparsity of SpargeAttn, MInference, and FlexPrefill is 0.5, 0.5, and 0.54.